zero product property
2w^2=16w

Answer by stanbon(75887)

zero product property
2w^2=16w
---
2w^2-16w = 0
Factor:
2w(w-8) = 0
----
Since the product is zero, one of the factors must be zero.
2w = 0 or w-8 = 0
w = 0 or w = 8
====================
Cheers,
Stan H.